# Enhancing Sample Selection Against Label Noise by Cutting Mislabeled Easy Examples

### v2

@@ Abstract
Sample selection is a prevalent approach in learning with noisy labels, aiming to identify confident samples for training. Although existing sample selection methods have achieved decent results by reducing the noise rate of the selected subset, they often overlook that not all mislabeled examples harm the model’s performance equally. In this paper, we demonstrate that mislabeled examples correctly predicted by the model early in the training process are particularly harmful to model performance. We refer to these examples as Mislabeled Easy Examples (MEEs). To address this, we propose Early Cutting, which introduces a recalibration step that employs the model’s later training state to re-select the confident subset identified early in training, thereby avoiding misleading confidence from early learning and effectively filtering out MEEs. Experiments on the CIFAR, WebVision, and full ImageNet-1k datasets demonstrate that our method effectively improves sample selection and model performance by reducing MEEs.
@@ 1 Introduction
Deep Neural Networks (DNNs) have achieved remarkable success, while heavily relies on the availability of high-quality, accurately annotated data (Han et al., 2020). In practice, collecting large-scale datasets with precise labels is challenging due to the high costs involved and the inherent subjectivity of manual annotation processes. Consequently, datasets often contain noisy labels, which can degrade the generalization performance of DNNs—a problem known as learning with noisy labels (LNL) (Natarajan et al., 2013). One prevalent approach to address LNL is sample selection, which aims to identify confident samples for training while discarding potentially mislabeled ones.
Sample selection methods can be categorized into two types: loss-based and dynamics-based. Loss-based methods rely on the assumption that clean samples tend to have smaller loss values than mislabeled samples (Han et al., 2018; Liu et al., 2020; Xia et al., 2021; Li et al., 2024a). In contrast, dynamics-based methods exploit the memorization effect of DNNs, which suggests that DNNs learn simple patterns first and then gradually fit the assigned label for each particular minority instance, including mislabeled samples (Liu et al., 2020; Zhang et al., 2021; Yuan et al., 2024). By analyzing the learning dynamics of DNNs, these methods aim to identify clean samples that are learned early and consistently throughout the training process (Yu et al., 2019; Xia et al., 2020a; Bai and Liu, 2021; Wei et al., 2022), considering them as confident samples for training. In recent years, dynamics-based methods have gained attention due to their ability to select Clean Hard Examples (CHEs)—challenging clean samples that are difficult to identify but crucial for achieving near-optimal generalization performance (Feldman and Zhang, 2020; Bai and Liu, 2021; Yuan et al., 2023).
Figure 1: (a) Test accuracy curves when the originally clean training subset is augmented with 4000 Mislabeled Easy Examples versus 4000 Mislabeled Hard Examples. Adding Mislabeled Easy Examples leads to a larger decrease in the model’s generalization performance. (b) Histogram illustrating the distribution of ImageNet-1k examples with 40% symmetric label noise, showing the epoch at which each example is first correctly predicted by the model during training. The horizontal axis represents the epoch when examples are first correctly predicted, and the vertical axis represents the number of examples predicted correctly at each epoch.
@@
Contributions. (1). We discover that mislabeled samples correctly predicted by the model early in training disproportionately harm model’s performance; we define these samples as Mislabeled Easy Examples (MEEs). MEEs are closer to the centers of their mislabeled classes in the feature space of models in early training stages, causing the model to easily learn incorrect patterns, which explains their especially harmful effect. (2). We introduce Early Cutting method, which recalibrates the confident data subset identified early in training by utilizing the model from later stages—a counterintuitive approach, given that later-stage models are typically regarded as less trustworthy.
@@
Sample Selection has been used in learning with noisy labels to improve the robustness of model training by prioritizing confident samples. An in-depth understanding of deep learning models, particularly their learning dynamics, has facilitated research in this area. Extensive studies on the Learning Dynamics of DNNs have revealed that difficult clean examples are typically learned in the later stages of training (Arpit et al., 2017; Yuan et al., 2024). This insight has led to training-time metrics that quantify sample “hardness”, such as learning speed (Jiang et al., 2021). These metrics inspire methods that leverage learning dynamics to select clean samples (Zhou et al., 2021; Maini et al., 2022). Various forms of Hard Label Noise have been studied, including asymmetric noise (Scott et al., 2013), instance-dependent noise (Xia et al., 2020b), natural noise (Wei et al., 2021b), adversarially crafted labels (Zhang et al., 2024a), open-set noise (Wei et al., 2021a), and subclass-dominant noise (Bai et al., 2023). These noise are designed from the perspective of the labels, aiming to simulate challenging real-world scenarios or malicious attacks. In contrast to prior studies that mainly focus on different types of label noise, our work offers a fresh perspective by re-examining sample selection methods that rely on a model’s early learning stages. We demonstrate that some samples hidden among those considered “confident” are, in fact, the most harmful. This contributes new insights into effectively identifying and handling mislabeled data.
@@ 2.1 Effects on Generalization from Mislabeled Examples Learned at Different Stages
Previous studies have shown that DNNs typically exhibit a specific learning pattern: they tend to learn simple and clean patterns first and gradually memorize more complex or mislabeled examples later (Arpit et al., 2017; Toneva et al., 2018). Based on this, some sample selection methods (Liu et al., 2020; Bai and Liu, 2021) trust the samples learned early by the model, treating them as high-probability clean samples. In our study, to distinguish the order in which the model learns different mislabeled examples, we refer to the definition in Yuan et al. (2023). Specifically, we consider that the model has learned a sample $(\mathbf{x}_{i},y_{i})$ at time $T_{i}$ if it consistently predicts the given label $y_{i}$ for both epoch $T_{i}-1$ and $T_{i}$, regardless of whether the label is correct. Formally, we define the learning time $LT_{i}$ of a sample $(\mathbf{x}_{i},y_{i})$ as:
$LT_{i}=\min\{T_{i}\mid\hat{y}_{i}^{T_{i}-1}=\hat{y}_{i}^{T_{i}}=y_{i}\},$ (1)
@@
where $\hat{y}_{i}^{t}$ denotes the model’s predicted label for instance $\mathbf{x}_{i}$ at epoch $t$. By tracking each sample’s learning time $LT_{i}$, we can analyze the order in which the model learns different samples and evaluate their impact on performance.
(b) Impact on model performance from mislabeled examples learned at different stages, using CIFAR-100.
Figure 2: Impact of mislabeled samples learned at different stages on model generalization performance. Subfigure 2(a) shows the scenario in the CIFAR-10 dataset, which contains 20,000 mislabeled samples (40% instance-dependent label noise) and 30,000 clean samples. We divided the 20,000 mislabeled samples into five groups based on the order in which an initial model learned them—from earliest to latest (ranging from $(0:20,000]$). Each group was combined with the 30,000 clean samples, creating datasets with approximately 12% label noise ($4,000/34,000$). New models were then trained on these datasets. As shown by the decreasing test accuracy, models trained on datasets containing earlier-learned mislabeled samples (e.g., “Clean +(0:4000]+(0:4000]+ ( 0 : 4000 ] Mislabeled”) exhibited lower generalization performance. Subfigure 2(b) shows similar findings on CIFAR-100.
@@
To investigate how the learning order of mislabeled examples affects their impact on the model’s generalization performance, we conduct experiments on the CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009) with 40% instance-dependent label noise (Xia et al., 2020b), resulting in 20,000 mislabeled eamples. We train an initial model $M_{i}$ on this noisy dataset and record the learning time $LT_{i}$ for each sample. We rank the mislabeled examples in ascending order of their learning times $LT_{i}$ and divide them into five groups of 4,000 mislabeled samples each, corresponding to learning stages from early to late: (0: 4000]delimited-(]:04000(0\,:\,4000]( 0 : 4000 ], (4000: 8000]delimited-(]:40008000(4000\,:\,8000]( 4000 : 8000 ], (8000: 12000]delimited-(]:800012000(8000\,:\,12000]( 8000 : 12000 ], (12000: 16000]delimited-(]:1200016000(12000\,:\,16000]( 12000 : 16000 ], and (16000: 20000]delimited-(]:1600020000(16000\,:\,20000]( 16000 : 20000 ]. For each group, we combine the 4,000 mislabeled examples with 30,000 clean examples to create new training datasets. This yields five datasets containing the same clean examples but different subsets of mislabeled examples grouped by their learning times $LT_{i}$.
@@
We then train new models from scratch on these datasets using the same architecture and hyperparameters, and evaluate their generalization performance on the clean test set. As shown in Figure 2, models trained on datasets with earlier-learned mislabeled examples (e.g., combining the earliest (0: 4000]delimited-(]:04000(0\,:\,4000]( 0 : 4000 ] mislabeled examples with 30,000 clean examples) exhibit significant lower generalization performance than those trained on datasets with later-learned mislabeled examples (e.g., combining the latest (16000: 20000]delimited-(]:1600020000(16000\,:\,20000]( 16000 : 20000 ] mislabeled examples with 30,000 clean examples). This indicates that mislabeled examples learned earlier by the model cause greater harm to generalization.
To validate this observation and minimize the impact of training stage differences on the learning order of mislabeled examples, we conducted additional experiments using a model pretrained on the clean CIFAR dataset, excluding the 20,000 mislabeled examples. As shown in Figure 3, whether starting from random initialization or using a pretrained model, the earlier-learned mislabeled examples are consistently learned faster by the model. This result further confirms our observation: mislabeled samples which are correctly predicted by the model early in the training process disproportionately degrade model performance.
@@ 2.2 Mislabeled Easy Examples
(b) The speed at which pretrained models on CIFAR-100 learn mislabeled examples from different stages.
Figure 3: Comparison of how pretrained models learn mislabeled examples from different learning stages. Subfigure 3(a) shows results on CIFAR-10 with 40% noise. We divided the mislabeled examples into five groups based on the order the initial model learned them, mixing each group with 30,000 clean examples to form datasets with approximately $12\%$ label noise ($4000/34000$). A model was pretrained on the 30,000 clean examples and then training on these new noisy datasets. Reference lines indicate the number of epochs required for the pretrained model to learn different sets of 2,000 mislabeled examples. The results reveal that earlier-learned mislabeled examples are also learned more quickly by the robust model. Subfigure 3(b) shows similar findings on CIFAR-100.
@@
As shown in the Figure 4(a), we visualize the mislabeled examples that are correctly predicted by the early-stage model using t-SNE (Van der Maaten and Hinton, 2008) in the feature space. Further, to quantify the model’s representations of mislabeled samples during early training, we compute the Euclidean distances from each mislabeled example learned by the early-stage model to the center of its true class and the center of its mislabeled class in the embedded feature space. We denote these two distances as $d_{\text{true}}$ and $d_{\text{mislabeled}}$, respectively. We then define the distance ratio $r=d_{\text{mislabeled}}/d_{\text{true}}$. If $r<1$, the example is closer to the mislabeled class center than to its true class center. As shown in the bottom row, MEEs exhibit a notably smaller median distance ratio ($0.830$), with more than half ($53.8\%$) of them having $r<1$. In contrast, the remaining mislabeled samples (non-MEEs) have a median ratio of $3.923$, and only $5.4\%$ are closer to the incorrect class.
@@
This phenomenon explains why MEEs are learned earlier: their misleading features align with the simple patterns of their given (incorrect) label classes that the model is tend to learn during the initial training stages. Thus, the early learning of MEEs has a disproportionately negative impact on the model’s generalization performance: since the model incorporates incorrect patterns associated with MEEs from the beginning, it disrupts the initial formation of simple and accurate feature representations. The erroneous features learned from MEEs become intertwined with the representations of clean data, making it challenging for the model to disentangle the clean patterns.
@@ 3 Methodology
Based on the analysis above, MEEs—mislabeled samples that the model learns easily during early training stages—can have a disproportionately negative impact on model generalization performance. Previous methods (Liu et al., 2020; Bai and Liu, 2021) often rely on trusting the model’s early learning stages or focusing on samples with small loss, are ineffective at filtering out MEEs due to their deceptive nature. To mitigate the influence of MEEs, we propose a novel sample selection strategy called Early Cutting. Early Cutting leverages the model’s state at a later training phase—specifically, at the early stopping epoch $t$ when the model begins to overfit—to re-evaluate and refine the subset of samples initially identified as confidently learned. The initial confident subset, $\mathcal{D}^{s}$, is formed by selecting samples with small learning times $LT_{i}$ (as defined in Eq. (1)). This captures examples that are learned quickly, encompassing both genuinely clean, easy samples and, crucially, the MEEs.
From this initially selected confident subset $\mathcal{D}^{s}$, our objective is to identify and subsequently remove samples that we characterize as MEEs in Section 2.2. These are samples that, when evaluated by the model $f_{\theta^{t}}$ at the early stopping epoch $t$, exhibit predictions that differ from their given noisy labels $y_{i})$, yet are made with high confidence and possess stable gradients.
Formally (algorithm for our method is provided in Appendix B.6), consider $\mathcal{D}^{s}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents input features, and $y_{i}\in\{1,2,\dots,K\}$ denotes the corresponding labels from $K$ classes. Let $f_{\theta^{t}}(\mathbf{x})$ be a model parameterized by $\theta^{t}$ at the early stopping epoch $t$, generating class probabilities via the softmax function:
@@
where $p_{i}^{(k)}=\Pr(y=k\,|\,\mathbf{x}_{i};\theta^{t})$ and $\sum_{k=1}^{K}p_{i}^{(k)}=1$. The predicted label $\hat{y}_{i}$ and the prediction confidence $c_{i}$ are given by: $\hat{y}_{i}=\arg\max_{k}\,p_{i}^{(k)}$, and $c_{i}=p_{i}^{(\hat{y}_{i})}$. The cross-entropy loss for sample $(\mathbf{x}_{i},y_{i})$ is: $L_{i}=-\log p_{i}^{(y_{i})}$.
@@
where $\delta$ and $\tau$ are thresholds for the loss and confidence, respectively. The rationale is that a high loss $L_{i}$ indicates that the model’s prediction at epoch $t$ disagrees with the given label $y_{i}$, and a high confidence $c_{i}$ implies that the model is very certain about its (contradictory) prediction $\hat{y}_{i}\neq y_{i}$. Therefore, samples satisfying both conditions are likely to be mislabeled, even if they were learned early. Relying solely on loss and confidence may not be sufficient, as some hard-to-learn samples may also exhibit high loss and high confidence due to their intrinsic difficulty. To further refine our selection, we introduce the concept of gradient stability. We compute the Euclidean norm of the gradient of the loss $L_{i}$ with respect to the input $\mathbf{x}_{i}$:
@@
A small gradient norm $g_{i}$ indicates that the loss $L_{i}$ is insensitive to small perturbations in $\mathbf{x}_{i}$, suggesting a strong (but potentially incorrect) association between the input features and the predicted label. MEEs tend to have low gradient norms because the model has confidently mislearned them, making the loss stable even under input perturbations. We refine $\mathcal{S}$ by selecting samples with high gradient stability:
@@
$\mathcal{S}^{\prime}=\left\{i\in\mathcal{S}\,\Big{|}\,g_{i}<\epsilon\right\},$ (5)
@@
where $\epsilon$ is a small threshold. This identification, is practically implemented by selecting samples whose metrics fall into predefined percentiles derived from their distributions within $\mathcal{D}^{s}$. Specifically, for all settings, we target the top $10\%$ for loss, top $20\%$ for confidence, and bottom $20\%$ for gradient norm. These percentile-based selections are intentionally kept fixed across all experimental settings to underscore the method’s general applicability and robustness, thereby obviating the need for dataset-specific hyperparameter tuning. Samples meeting these criteria are classified as MEEs and subsequently removed from $\mathcal{D}^{s}$, yielding a refined subset for further training.
@@
Notably, the proposed method operates on $\mathcal{D}^{s}$, the subset of samples learned quickly during the initial training phase. Such early-learned subsets are known to often contain significant redundancy, with multiple examples representing similar, dominant data patterns (Feldman, 2020; Feldman and Zhang, 2020; Yuan et al., 2024). This inherent redundancy contributes to the robustness of our MEEs removal strategy. Firstly, it provides resilience against the inadvertent removal of a small number of clean samples, as their informational content is likely preserved by other remaining examples. Secondly, this characteristic makes the outcome less sensitive to the precise percentile thresholds used for MEEs selection. A sensitivity analysis of these thresholds is presented in Section 4.3, an ablation study is shown in Appendix B.9, and it transferability is shown in Appendix B.10.

### v3

@@ Abstract
Sample selection is a prevalent approach in learning with noisy labels, aiming to identify confident samples for training. Although existing sample selection methods have achieved decent results by reducing the noise rate of the selected subset, they often overlook that not all mislabeled examples harm the model’s performance equally. In this paper, we demonstrate that mislabeled examples correctly predicted by the model early in the training process are particularly harmful to model performance. We refer to these examples as Mislabeled Easy Examples (MEEs). To address this, we propose Early Cutting, which introduces a recalibration step that employs the model’s later training state to re-select the confident subset identified early in training, thereby avoiding misleading confidence from early learning and effectively filtering out MEEs. Experiments on the CIFAR, WebVision, and full ImageNet-1k datasets demonstrate that our method effectively improves sample selection and model performance by reducing MEEs. Our implementation can be found at https://github.com/tmllab/2025_NeurIPS_MEE.
@@ 1 Introduction
Deep Neural Networks (DNNs) have achieved remarkable success, while heavily relying on the availability of high-quality, accurately annotated data. In practice, collecting large-scale datasets with precise labels is challenging due to the high costs involved and the inherent subjectivity of manual annotation processes. Consequently, datasets often contain noisy labels, which can degrade the generalization performance of DNNs—a problem known as learning with noisy labels (LNL) (Natarajan et al., 2013). One prevalent approach to address LNL is sample selection, which aims to identify confident samples for training while discarding potentially mislabeled ones.
Sample selection methods can be categorized into two types: loss-based and dynamics-based. Loss-based methods rely on the assumption that clean samples tend to have smaller loss values than mislabeled samples (Han et al., 2018; Liu et al., 2020; Xia et al., 2021; Li et al., 2024). In contrast, dynamics-based methods exploit the memorization effect of DNNs, which suggests that DNNs learn simple patterns first and then gradually fit the assigned label for each particular minority instance, including mislabeled samples (Liu et al., 2020; Zhang et al., 2021; Yuan et al., 2024). By analyzing the learning dynamics of DNNs, these methods aim to identify clean samples that are learned early and consistently throughout the training process (Yu et al., 2019; Xia et al., 2020a; Bai and Liu, 2021; Wei et al., 2022), considering them as confident samples for training. In recent years, dynamics-based methods have gained attention due to their ability to select Clean Hard Examples (CHEs)—challenging clean samples that are difficult to identify but crucial for achieving near-optimal generalization performance (Feldman and Zhang, 2020; Bai and Liu, 2021; Yuan et al., 2023).
Figure 1: (a) Test accuracy curves when the originally clean training subset is augmented with 4000 Mislabeled Easy Examples versus 4000 Mislabeled Hard Examples (see Section 2.1 for setup). Adding Mislabeled Easy Examples leads to a larger decrease in the model’s generalization performance. (b) Histogram illustrating the distribution of ImageNet-1k examples with 40% symmetric label noise, showing the epoch at which each example is first correctly predicted by the model during training. The horizontal axis represents the epoch when examples are first correctly predicted, and the vertical axis represents the number of examples predicted correctly at each epoch.
@@
Our main contributions can be summarized as follows:
  • 1.
    We discover that mislabeled samples correctly predicted by the model early in training disproportionately harm model’s performance; we define these samples as Mislabeled Easy Examples (MEEs). MEEs are closer to the centers of their mislabeled classes in the feature space of models in early training stages, causing the model to easily learn incorrect patterns.
  • 2.
    We introduce Early Cutting method, which recalibrates the confident data subset identified early in training by utilizing the model from later stages—a counterintuitive approach, given that later-stage models are typically regarded as less trustworthy.
@@
Sample Selection has been used in learning with noisy labels to improve the robustness of model training by prioritizing confident samples. An in-depth understanding of deep learning models, particularly their learning dynamics, has facilitated research in this area. Extensive studies on the Learning Dynamics of DNNs have revealed that difficult clean examples are typically learned in the later stages of training (Arpit et al., 2017; Yuan et al., 2024, 2025). This insight has led to training-time metrics that quantify sample “hardness”, such as learning speed (Jiang et al., 2021). These metrics inspire methods that leverage learning dynamics to select clean samples (Zhou et al., 2021; Maini et al., 2022). Various forms of Hard Label Noise have been studied, including asymmetric noise (Scott et al., 2013), instance-dependent noise (Xia et al., 2020b), natural noise (Wei et al., 2021b), adversarially crafted labels (Zhang et al., 2024a), open-set noise (Wei et al., 2021a), and subclass-dominant noise (Bai et al., 2023). These noise types are designed from the perspective of the labels, aiming to simulate challenging real-world scenarios or malicious attacks. In contrast to prior studies that mainly focus on different types of label noise, our work offers a fresh perspective by re-examining sample selection methods that rely on a model’s early learning stages. We demonstrate that some samples hidden among those considered “confident” are, in fact, the most harmful. This contributes new insights into effectively identifying and handling mislabeled data.
@@ 2 Our Observations
Figure 2: Impact of mislabeled samples learned at different stages on model generalization performance. Subfigure 2 shows the scenario in the CIFAR-10 dataset, which contains 20,000 mislabeled samples (40% instance-dependent label noise) and 30,000 clean samples. We divided the 20,000 mislabeled samples into five groups based on the order in which an initial model learned them—from earliest to latest (ranging from $(0:20,000]$). Each group was combined with the 30,000 clean samples, creating datasets with approximately 12% label noise ($4,000/34,000$). New models were then trained on these datasets. As shown by the decreasing test accuracy, models trained on datasets containing earlier-learned mislabeled samples (e.g., “Clean +(0:4000]+(0:4000] Mislabeled”) exhibited lower generalization performance. Subfigure 6 shows similar findings on CIFAR-100.
@@ 2.1 Effects on Generalization from Mislabeled Examples Learned at Different Stages
Previous studies have shown that DNNs typically exhibit a specific learning pattern: they tend to learn simple and clean patterns first and gradually memorize more complex or mislabeled examples later (Arpit et al., 2017; Toneva et al., 2018). Based on this, some sample selection methods (Liu et al., 2020; Bai and Liu, 2021) trust the samples learned early by the model, treating them as high-probability clean samples. In our study, to distinguish the order in which the model learns different mislabeled examples, we refer to the definition in Yuan et al. (2023). Specifically, we consider that the model has learned a sample $(\mathbf{x}_{i},\tilde{y}_{i})$ at time $E_{i}$ if it consistently predicts the given label $\tilde{y}_{i}$ for both epoch $E_{i}-1$ and $E_{i}$, regardless of whether the label is correct. Formally, we define the learning time $LT_{i}$ of a sample $(\mathbf{x}_{i},\tilde{y}_{i})$ as:
$LT_{i}=\min\{E_{i}\mid\hat{y}_{i}^{E_{i}-1}=\hat{y}_{i}^{E_{i}}=\tilde{y}_{i}\},$ (1)
@@
where $\hat{y}_{i}^{t}$ denotes the model’s predicted label for instance $\mathbf{x}_{i}$ at epoch $e$. By tracking each sample’s learning time $LT_{i}$, we can analyze the order in which the model learns different samples and evaluate their impact on performance.
@@
To investigate how the learning order of mislabeled examples affects generalization, we conducted experiments on CIFAR-10 and CIFAR-100 with 40% instance-dependent label noise. First, we trained an initial model on the noisy dataset to record the learning time $LT_{i}$ for each sample. Based on these times, we partitioned the 20,000 mislabeled examples into five sequential groups of 4,000, from earliest-learned to latest-learned. Each group was then combined with 30,000 clean examples to form five distinct training datasets.
@@
We then trained new models from scratch on these datasets. As shown in Figure 2, the results are unambiguous: models trained with the earliest-learned group of mislabeled examples exhibit significantly lower generalization performance than those trained with later-learned groups. This clearly indicates that mislabeled examples learned earlier by a model cause greater harm to its generalization. To validate this observation, we repeated the experiment using a model pretrained on only the clean data. As shown in Figure 3, the model still learns the MEEs from the “earliest” group much faster than those from later groups. This confirms that these samples are inherently easy for the model to learn, regardless of the training starting point.
@@ 2.2 Mislabeled Easy Examples
Figure 3: Comparison of how pretrained models learn mislabeled examples from different learning stages. Subfigure 3 shows results on CIFAR-10 with 40% noise. We divided the mislabeled examples into five groups based on the order the initial model learned them, mixing each group with 30,000 clean examples to form datasets with approximately $12\%$ label noise ($4000/34000$). A model was pretrained on the 30,000 clean examples and then trained on these new noisy datasets. Reference lines indicate the number of epochs required for the pretrained model to learn different sets of 2,000 mislabeled examples. The results reveal that earlier-learned mislabeled examples are also learned more quickly by the robust model. Subfigure 7 shows similar findings on CIFAR-100.
@@
As shown in the Figure 4(a), we visualize the mislabeled examples that are correctly predicted by the early-stage model using t-SNE (Van der Maaten and Hinton, 2008) in the feature space. For a detailed visualization of how this feature space and the associated distance ratios evolve at later training stages, please see Appendix E. Further, to quantify the model’s representations of mislabeled samples during early training, we compute the Euclidean distances from each mislabeled example learned by the early-stage model to the center of its true class and the center of its mislabeled class in the embedded feature space. We denote these two distances as $d_{\text{true}}$ and $d_{\text{mislabeled}}$, respectively. We then define the distance ratio $r=d_{\text{mislabeled}}/d_{\text{true}}$. If $r<1$, the example is closer to the mislabeled class center than to its true class center. As shown in the bottom row, MEEs exhibit a notably smaller median distance ratio ($0.830$), with more than half ($53.8\%$) of them having $r<1$. In contrast, the remaining mislabeled samples (non-MEEs) have a median ratio of $3.923$, and only $5.4\%$ are closer to the incorrect class.
@@
This phenomenon explains why MEEs are learned earlier: their misleading features align with the simple patterns of their given (incorrect) label classes that the model is tend to learn during the initial training stages. Thus, the early learning of MEEs has a disproportionately negative impact on the model’s generalization performance: since the model incorporates incorrect patterns associated with MEEs from the beginning, it disrupts the initial formation of simple and accurate feature representations. The erroneous features learned from MEEs become intertwined with the representations of clean data, making it challenging for the model to disentangle the clean patterns. To further quantify this detrimental impact at a microscopic level, we conducted a supplementary analysis using influence functions (Koh and Liang, 2017) to measure the precise impact of individual samples on model generalization. This analysis, detailed in Appendix B, provides direct quantitative evidence that MEEs are substantially more harmful to the model’s performance on clean data than other mislabeled examples.
@@ 3 Methodology
Based on the analysis above, MEEs, mislabeled samples that the model learns easily during early training stages, can have a disproportionately negative impact on model generalization performance. Previous methods (Liu et al., 2020; Bai and Liu, 2021) often rely on trusting the model’s early learning stages or focusing on samples with small loss, are ineffective at filtering out MEEs due to their deceptive nature. To mitigate the influence of MEEs, we propose a novel sample selection strategy called Early Cutting. Early Cutting leverages the model’s state at a later training phase—specifically, at the early stopping epoch $t$ when the model begins to overfit—to re-evaluate and refine the subset of samples initially identified as confidently learned. The initial confident subset, $\mathcal{D}^{s}$, is formed by selecting samples with small learning times $LT_{i}$ (as defined in Eq. (1)). This captures examples that are learned quickly, encompassing both genuinely clean, easy samples and, crucially, the MEEs.
From this initially selected confident subset $\mathcal{D}^{s}$, our objective is to identify and subsequently remove samples that we characterize as MEEs in Section 2.2. These are samples that, when evaluated by the model $f_{\theta^{t}}$ at the early stopping epoch $t$, exhibit predictions that differ from their given noisy labels $\tilde{y}_{i})$, yet are made with high confidence and possess stable gradients.
Formally, consider $\mathcal{D}^{s}={(\mathbf{x}_{i},\tilde{y}_{i})}_{i=1}^{N}$, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents input features, and $\tilde{y}_{i}\in\mathcal{Y}$ denotes the corresponding observed labels from $K$ classes. Let $f_{\theta^{t}}(\mathbf{x})$ be a model parameterized by $\theta^{t}$ at the early stopping epoch $t$, generating class probabilities via the softmax function:
@@
where $p_{i}^{(k)}$is the model’s output probability for class k. The predicted label $\hat{y}_{i}$ and the prediction confidence $c_{i}$ are given by: $\hat{y}_{i}=\arg\max_{k}\,p_{i}^{(k)}$, and $c_{i}=p_{i}^{(\hat{y}_{i})}$. The cross-entropy loss for sample $(\mathbf{x}_{i},\tilde{y}_{i})$ is: $L_{i}=-\log p_{i}^{(\tilde{y}_{i})}$.
@@
where $\delta$ and $\tau$ are thresholds for the loss and confidence, respectively. The rationale is that a high loss $L_{i}$ indicates that the model’s prediction at epoch $t$ disagrees with the given label $\tilde{y}_{i}$, and a high confidence $c_{i}$ implies that the model is very certain about its (contradictory) prediction $\hat{y}_{i}\neq\tilde{y}_{i}$. Therefore, samples satisfying both conditions are likely to be mislabeled, even if they were learned early. Our method remains robust even if the set $\mathcal{S}$ is empty, as the refinement step would simply be bypassed. Relying solely on loss and confidence may not be sufficient, as some hard-to-learn samples may also exhibit high loss and high confidence due to their intrinsic difficulty. To further refine our selection, we introduce the concept of gradient stability. We compute the Euclidean norm of the gradient of the loss $L_{i}$ with respect to the input $\mathbf{x}_{i}$:
@@
A small gradient norm $g_{i}$ indicates that the loss $L_{i}$ is insensitive to small perturbations in $\mathbf{x}_{i}$, suggesting a strong (but potentially incorrect) association between the input features and the predicted label. MEEs tend to have low gradient norms because the model has confidently mislearned them, making the loss stable even under input perturbations. We refine $\mathcal{S}$ by selecting samples with high gradient stability (where $\epsilon$ is a threshold):
$\mathcal{S}^{\prime}=\left\{i\in\mathcal{S}\,\Big|\,g_{i}<\epsilon\right\}.$ (5)
Operational Definition 1 (Mislabeled Easy Examples (MEEs)).
Operationally, from the set of early-learned samples $\mathcal{D}^{s}$, we define the subset of Mislabeled Easy Examples (MEEs) as those samples $i$ that satisfy the conditions of high loss, high confidence, and low input gradient norm. Formally:
@@
$\text{MEEs}=\left\{i\in\mathcal{D}^{s}\mid(L_{i}>\delta)\land(c_{i}>\tau)\land(g_{i}<\epsilon)\right\}.$ (6)
@@
This identification is practically implemented by selecting samples whose metrics fall into predefined percentiles derived from their distributions within $\mathcal{D}^{s}$. Specifically, for all settings, we target the top $10\%$ for loss, top $20\%$ for confidence, and bottom $20\%$ for gradient norm, which were determined on a validation set. Additionally, we set the early cutting rate to 1.5. Samples meeting these criteria are classified as MEEs and subsequently removed from $\mathcal{D}^{s}$, yielding a refined subset for further training. We formalize the complete Early Cutting procedure in Algorithm 1.
Algorithm 1 Iterative Sample Selection with Early Cutting
1:Training data $\mathcal{D}^{0}$; Number of iterations $I_{\text{rate}}$; Early cutting rate $\gamma$; Thresholds $\delta,\tau,\epsilon$
2:Trained model parameters $\theta^{*}$
3:for $i=1$ to $I_{\text{rate}}$ do
4:  1. Base Sample Selection:
5:  Compute learning times $LT_{j}$ for all $(\mathbf{x}_{j},\tilde{y}_{j})\in\mathcal{D}^{i-1}$ using Eq. (1)
6:  Select the initial early-learned subset $\mathcal{D}^{s}$ based on smallest learning times.
7:  2. Early Cutting:
8:  Select candidate subset $\mathcal{D^{\prime}}^{s}$ from $\mathcal{D}^{s}$ (e.g., the $\frac{1}{\gamma}$ proportion with the earliest learning times).
9:  Compute loss $L_{j}$, confidence $c_{j}$, and gradient norm $g_{j}$ for all samples in $\mathcal{D^{\prime}}^{s}$.
10:  Identify the set of MEEs in $\mathcal{D^{\prime}}^{s}$ according to Definition 1:
11:   $\text{MEEs}=\left\{j\in\mathcal{D^{\prime}}^{s}\mid(L_{j}>\delta)\land(c_{j}>\tau)\land(g_{j}<\epsilon)\right\}$
12:  Create the refined subset by removing the identified MEEs: $\mathcal{D}^{s}_{\text{refined}}\leftarrow\mathcal{D}^{s}\setminus\text{MEEs}$.
13:  Update the training data for the next iteration: $\mathcal{D}^{i}\leftarrow\mathcal{D}^{s}_{\text{refined}}$.
14:end for
15:Final Training Phase: Train a model from scratch on the final refined set $\mathcal{D}^{I_{\text{rate}}}$ until convergence.
16:return Trained model parameters $\theta^{*}$.
@@
Notably, the proposed method operates on $\mathcal{D}^{s}$, the subset of samples learned quickly during the initial training phase. Such early-learned subsets are known to often contain significant redundancy, with multiple examples representing similar, dominant data patterns (Feldman, 2020; Feldman and Zhang, 2020; Yuan et al., 2024). This inherent redundancy contributes to the robustness of our MEEs removal strategy. Firstly, it provides resilience against the inadvertent removal of a small number of clean samples, as their informational content is likely preserved by other remaining examples. Secondly, this characteristic makes the outcome less sensitive to the precise percentile thresholds used for MEEs selection. A sensitivity analysis of these thresholds is presented in Section 4.3, an ablation study is shown in Appendix D.8, and it transferability is shown in Appendix D.9.